OpenStudy (julyahx1):

The two triangles below are similar. What is the similarity ratio of ΔXYZ to ΔABC? Triangle XYZ is shown with XZ measuring 10 and YZ measuring 30. Triangle ABC is shown with side AC measuring 15. 1:2 2:1 3:2 2:3 I'm thinking the answer is D but i am not 100% sure. Please someone help. Thank you!

OpenStudy (julyahx1):

@steve816 please help me

OpenStudy (julyahx1):

@sweetburger please help me

OpenStudy (idkanything):

@3mar

OpenStudy (3mar):

Well, I am here

OpenStudy (julyahx1):

please can u help me 3mar

OpenStudy (julyahx1):

i think its D but i don't know 100%

OpenStudy (3mar):

Of course You need it now?

OpenStudy (julyahx1):

yes

OpenStudy (3mar):

Ok

OpenStudy (idkanything):

ya idk it might be D

OpenStudy (3mar):

I am sure it is D. Correct JULYAHX1 because the ratio of XZ/AC=10/15=2/3

OpenStudy (julyahx1):

oh ok thanks so much for the clarification!

OpenStudy (3mar):

So welcome! Any help any time

OpenStudy (3mar):

Thanks for the medal!

OpenStudy (julyahx1):

i have just one more question i need for a clarification if you're not busy? and yeah no problem (:

OpenStudy (julyahx1):

u there?

OpenStudy (3mar):

ok